# T3: A Novel Zero-shot Transfer Learning Framework Iteratively Training on an Assistant Task for a Target Task

### v1

@@ 1 Introduction
In recent years, long text summarization has gradually played an important role in various real-world scenarios such as current affairs commentary, book summaries, news aggregation, and literature research, benefiting from the fast developing of Large Language Models (LLMs) Nallapati et al. (2016); Vaswani et al. (2023); Dubey et al. (2024); Islam and Moushi (2024). Different from classical summarization tasks, which typically summarize texts of around 500 words or less into 100 or fewer words, long text summarization involves dealing with much longer texts. Conceivably, as the original text length increases, the amount of contextual details and relevance required in the summary also grows, bringing new challenges to LLMs.
@@ 3 Methodology
In this section, we first demonstrate the general workflow of proposed T3, and then formulate the summarization task under T3, thereafter dive into the detailed training and test processes.
3.1 T3 General Workflow
The general workflow of T3 is shown in Figure 1. The keypoint of T3 is that: in the whole training process, both the assistant task and the target task share the samples of the assistant task. At first, the used LLM  [^1] GPT-4o is recommended, while any other large models with sufficient parameter capacity to possess emergence could be considered. acquires experience to deal with the assistant task through the labeled data, which means a deeper comprehension of the samples. Subsequently, LLM iteratively tries to accomplish the target task on these same samples and gain the experience for the target task by leveraging the previous learned contextual experiences, as well as being modulated by the predefined conditions. Finally, the model applies the experiences required for both the assistant and target tasks during the test process.
3.2 T3 for Summarization
Now the general workflow of T3 is implemented to our target task: summarization, with the question-answering issue as the assistant task.
Figure 2: Training and test process of T3 for summarization task.
3.2.1 Task Formulation
@@
Figure 3: Test process of T3 for summarization task on QA dataset.
3.2.2 Training Process
@@
The overview of T3 for summarization is illustrated in Figure 2. The training process can be found at the top of the figure, while the test process is demonstrated at the bottom. Both inputs and outputs in these processes are written in bold in the figure. Furthermore, the detailed information of the training process is given in Algorithm 1 in Appendix A.1, including the following key steps:
Updating $Exp_{QA}$: At first, both summary and QA generation experiences $Exp_{Sum}$ and $Exp_{QA}$ are initialized as $null$. Then, LLM generates new QA pairs $QA_{i}^{\prime}$ for each text $D_{i}$ and updates $Exp_{QA}$ according to several rules concluded by itself through comparing $QA_{i}\textquoteright$ ’ with the ground truth $QA_{i}$. The prompt $P_{QA}$ is predefined to control all steps mentioned in this part.
Updating $Exp_{Sum}$: In this part, LLM follows the predefined prompts $P_{Sum}$ and try to summarize the text $D_{i}$. As an prior-knowledge that might bring extra contextual information, both $QA_{i}$ and $QA_{i}\textquoteright$ ’ are also taken into account as the inputs. After the current summary $Sum_{i}^{\prime}$ is generated, three thresholds are used to investigate whether the generated content is well enough. LLM would repeat the summary generation step and update its experience $Exp_{Sum}$ every iteration, till the stop conditions are fulfilled.
Thresholds Used in Iteration: The iterative process mentioned above continues until the generated summary $Sum_{i}^{\prime}$ simultaneously meets the required thresholds for similarity, readability Flesch (1948), and compression rate (or the number of iterations reach the preset maximum iteration times $K$). In details, readability R is calculated as:
$\text{C}=\text{length}(Summary)/\text{length}(Text)$ (2)
which is used to control the length of the generated summary.
3.2.3 Test Process on Summarization Dataset
According to the bottom of Figure 2, during the test process, LLM would first use $Exp_{QA}$ learned from the training process to generated auxiliary question-answer pairs, and bring them with the updated $Exp_{Sum}$ to generate the final summary. Prompts in test process are given in Appendix A.4.
3.2.4 Test Process on QA Dataset
4 Experiments
4.1 Settings
4.1.1 Datasets
4.1.2 Evaluation Metrics
FactSumm FactSum uses named-entity recognition and relation extraction to extract facts from the source text and the generated summary to evaluate fact consistency by comparing extracted facts represented as triples (subject, relation, object) Zhang et al. (2021).
ChatGPT evaluates factual inconsistency in summaries, outperforming previous methods with fewer hallucinations Luo et al. (2023).
4.1.3 Baseline Models
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
BBC summary
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.4019 | 0.4171 0.1814 | 0.2145 0.3697 | 0.3944 0.1086 | 0.1094 0.6364 | 0.7012
GPT-4-turbo 0.3733 | 0.3683 0.1476 | 0.1414 0.3373 | 0.3327 0.0851 | 0.0769 0.6091 | 0.6901
GPT-3.5-turbo 0.3936 | 0.3788 0.1775 | 0.1695 0.3604 | 0.3478 0.1050 | 0.0858 0.5752 | 0.4780
Gemini-1.5-pro 0.3362 | 0.3510 0.1148 | 0.1258 0.3036 | 0.3169 0.0505 | 0.0647 0.2588 | 0.6049
Gemini-1.0-pro 0.3490 | 0.3358 0.1280 | 0.1392 0.3175 | 0.3117 0.0599 | 0.0722 0.3576 | 0.5561
Claude-3.5-sonnet 0.3908 | 0.4354 0.1659 | 0.2113 0.3588 | 0.4036 0.0970 | 0.1300 0.6343 | 0.6583
Claude-3.0-sonnet 0.4123 | 0.4090 0.1826 | 0.1808 0.3778 | 0.3739 0.1108 | 0.1074 0.6002 | 0.5043
NarraSum
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.3114 | 0.2886 0.0731 | 0.0858 0.2844 | 0.2566 0.0268 | 0.0255 0.9775 | 0.9969
GPT-4-turbo 0.2914 | 0.2727 0.0772 | 0.0742 0.2544 | 0.2445 0.0336 | 0.0360 0.9775 | 0.9744
GPT-3.5-turbo 0.2860 | 0.2528 0.0752 | 0.0588 0.2650 | 0.2300 0.0279 | 0.0215 0.9520 | 0.9591
Gemini-1.5-pro 0.2599| 0.2695 0.0521 | 0.0539 0.2500 | 0.2532 0.0175 | 0.0199 0.9326 | 0.9357
Gemini-1.0-pro 0.2897 | 0.2825 0.0626 | 0.0574 0.2545 | 0.2636 0.0165 | 0.0209 0.9285 | 0.9122
Claude-3.5-sonnet 0.2919 | 0.2978 0.0717 | 0.0779 0.2673 | 0.2723 0.0310 | 0.0323 0.9704 | 0.9814
Claude-3.0-sonnet 0.2990 | 0.2587 0.0675 | 0.0547 0.2645 | 0.2251 0.0274 | 0.0205 0.9469 | 0.9561
NLQuAD
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.2437 | 0.3100 0.0848 | 0.1303 0.2271 | 0.2927 0.0033 | 0.0142 0.5981 | 0.6395
GPT-4-turbo 0.2419 | 0.2366 0.0739 | 0.0736 0.2216 | 0.2162 0.0044 | 0.0036 0.5333 | 0.4916
GPT-3.5-turbo 0.2150 | 0.2221 0.0775 | 0.0830 0.2007 | 0.2069 0.0027 | 0.0027 0.5530 | 0.4464
Gemini-1.5-pro 0.2299 | 0.2419 0.0697 | 0.0797 0.2134 | 0.2244 0.0025 | 0.0047 0.5085 | 0.6509
Gemini-1.0-pro 0.2626 | 0.2811 0.0929 | 0.1105 0.2455 | 0.2648 0.0054 | 0.0125 0.6457 | 0.7164
Claude-3.5-sonnet 0.3203 | 0.4017 0.1280 | 0.1874 0.3006 | 0.3792 0.0235 | 0.0500 0.6571 | 0.8276
Claude-3.0-sonnet 0.3235 | 0.3372 0.1270 | 0.1327 0.3016 | 0.3122 0.0197 | 0.0237 0.5889 | 0.6169
FairytaleQA
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.1334 | 0.1486 0.0315 | 0.0372 0.1234 | 0.1378 0.0004 | 0.0007 0.9855 | 0.9892
GPT-4-turbo 0.1468 | 0.1414 0.0351 | 0.0336 0.1345 | 0.1284 0.0009 | 0.0010 0.9744 | 0.9767
GPT-3.5-turbo 0.1222 | 0.1223 0.0298 | 0.0307 0.1131 | 0.1074 0.0003 | 0.0004 0.9982 | 0.9737
Gemini-1.5-pro 0.1540 | 0.1756 0.0380 | 0.0434 0.1430 | 0.1546 0.0010 | 0.0013 0.9723 | 0.9853
Gemini-1.0-pro 0.1414 | 0.1343 0.0325 | 0.0243 0.1422 | 0.1339 0.0003 | 0.0014 0.9784 | 0.9424
Claude-3.5-sonnet 0.1597 | 0.2096 0.0373 | 0.0607 0.1494 | 0.1952 0.0017 | 0.0047 0.9956 | 0.9997
Claude-3.0-sonnet 0.1720 | 0.1891 0.0408 | 0.0490 0.1581 | 0.1735 0.0017 | 0.0033 0.9747 | 0.9892
4.2 Implementation Details
The baseline summary of each text for each LLM is generated by using original text and the following prompt: "You are one helpful text assistant. Based on the input text, provide a summary in proper length". Correspondently, each LLM with T3 is guided by the experience and prompts derived from training process.
For BBC summary and NarraSum, the provided summary was used as the reference, while for FairytaleQA and NLQuAD datasets, the original text was used as the reference. As mentioned in Part 3.2.4, T3 can also deal with QA datasets for summarization task. When the dataset does not provide reference summaries, the generated summary can be directly compared to the original text through certain automated metrics tools (e.g., ROUGE) Louis and Nenkova (2009). According to Saggion and Poibeau (2013), the overlap of the generated summaries with the original text is important in assessing information coverage and the validity.
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Readbility Factscore
Claude-3.5-sonnet BBC summary
Full T3 0.4354 0.2113 0.4036 0.1300 54.71 0.6583
w/o summary experience 0.4026 0.1683 0.3599 0.1331 55.08 0.5832
w/o qa experience 0.4155 0.2072 0.4005 0.3190 55.64 0.5857
GPT-4o NLQuAD
Full T3 0.3100 0.1303 0.2927 0.0142 44.53 0.6395
w/o summary experience 0.3628 0.1529 0.3402 0.0309 42.57 0.7785
w/o qa experience 0.4116 0.2029 0.3899 0.0531 45.77 0.6523
Table 2: Ablation study results on BBC summary and NLQuAD datasets.
4.3 Results
Table 1 illustrates the performance of all used models with (right "T3" column) and without (left "w/o" column) T3 on all adopted datasets. The better one of the two values from each row of the two columns is marked in bold. Especially, for each row where all results from columns with T3 are better than those columns without T3, the row is colored blue. From this table, we observe that:
(1) LLMs using T3 outperforms themselves without T3 under most metrics of most datasets, which demonstates the feasibility and effectiveness of T3. Specially, the average improvement rates of GPT-4o, Gemini-1.5-pro, and Claude-3.5-sonnet using T3 compared to the baselines without T3 across different tasks are calculated and shown that these LLMs with T3 nearly improves the ROUGE score by 14%, the BLEU score by 35%, and the Factscore score by 16%.
4.4 Ablation Study
We further analyze the influence of different components of T3. Table 2 shows the experimental results for Claude-3.5-sonnet after removing summary experience and QA generation experience on the BBC summary dataset, and the results for GPT-4o after removing summary experience and QA generation experience on the NLQuAD dataset. For models with less reasoning ability, longer inputs and complex instructions of T3 might diminish their performance, while for more advanced models, each component of T3 could be utilized to enhance their overall effectiveness.
5 Conclusion
Limitations
Ethics Statement
Appendix A Appendix
A.1 Training Steps of T3 for Summarization
@@ A.2 Prompt for QA Generation and Experience Updating in Training Process
You are one helpful text assistant and you are good at learning new experience. First please help us to do the following:
1. Based on input text, provided questions and QA pair generation experience, generate questions and corresponding answers to original text that can reflect all the subjects and all the key points of input article. The format of the generated question pairs should be same as the provided question pairs.
2. Update the learning experience of how to generate better question pairs during the question generation process by comparing the generated questions and original question pairs based on the original text or other methods. You need to consider all the key points to make a good QA pairs and the learning experience should be general. DO NOT INCLUDE DETAIL EXAMPLES BUT PROVIDE rules AS DETAILED AS POSSIBLE.DO NOT EXCEED 600 WORDS.
@@
{ "Generated_QA_pairs":
{ "1": { "Question": "xxxx", "Answer": "xxxx"},
"2": { "Question": "xxxx", "Answer": "xxxx"},
"QA_generation_experience": "Answer point by point text in one line"}

### v3

@@ 1 Introduction
In recent years, long text summarization has gradually played an important role in various real-world scenarios such as current affairs commentary, book summaries, news aggregation, and literature research, benefiting from the fast development of Large Language Models (LLMs) Nallapati et al. (2016); Vaswani et al. (2023); Dubey et al. (2024); Islam and Moushi (2024). Unlike classical summarization tasks, which typically summarize texts of around 500 words or less into 100 or fewer words, long text summarization involves dealing with much longer texts. Clearly, as the original text length increases, the amount of contextual details and relevance required in the summary also grows, bringing new challenges to LLMs.
@@ 2.4 Zero-shot Transfer Learning
Figure 1: Training and test process of T3 for summarization task.
@@ 3 Methodology
In this section, we first formulate the summarization task under T3, and dive into the detailed training and test processes. Then, we conclude the general form of the T3 workflow.
3.1 Task Formulation
@@
3.2 Training Process
@@
The overview of T3 for summarization is illustrated in Figure 1. The training process can be found at the top of the figure, while the test process is demonstrated at the bottom. Both inputs and outputs in these processes are written in bold in the figure. Furthermore, the detailed information of the training process is given in Algorithm 1, including the following key steps:
@@
Step 1 Updating $Exp_{QA}$: At first, both summary and QA generation experiences $Exp_{Sum}$ and $Exp_{QA}$ are initialized as $null$. Then, LLM generates new QA pairs $QA_{i}^{\prime}$ for each text $D_{i}$ and updates $Exp_{QA}$ according to several rules concluded by itself through comparing $QA_{i}\textquoteright$ ’ with the ground truth $QA_{i}$. The prompt $P_{QA}$ is predefined to control all steps mentioned in this part.
Step 2 Updating $Exp_{Sum}$: In this part, LLM follows the predefined prompts $P_{Sum}$ and try to summarize the text $D_{i}$. As an prior-knowledge that might bring extra contextual information, both $QA_{i}$ and $QA_{i}\textquoteright$ ’ are also taken into account as the inputs. After the current summary $Sum_{i}^{\prime}$ is generated, three thresholds are used to investigate whether the generated content is well enough. LLM would repeat the summary generation step and update its experience $Exp_{Sum}$ every iteration, till the stop conditions are fulfilled.
Step 3 Thresholds Used in Iteration: The iterative process mentioned above continues until the generated summary $Sum_{i}^{\prime}$ simultaneously meets the required thresholds for similarity, readability Flesch (1948), and compression rate (or the number of iterations reach the preset maximum iteration times $K$). In details, readability R is calculated as:
$\text{C}=\text{length}(Summary)/\text{length}(Text)$ (2)
which is used to control the length of the generated summary.
3.3 Test Process on Summarization Dataset
According to the bottom of Figure 1, during the test process, LLM would first use $Exp_{QA}$ learned from the training process to generated auxiliary question-answer pairs, and bring them with the updated $Exp_{Sum}$ to generate the final summary. Prompts in test process are given in Appendix A.4 and a step-by-step use case is shown in Appendix A.10.3.
3.4 Test Process on QA Dataset
Figure 2: Test process of T3 for summarization task on QA dataset.
3.5 T3 General Workflow
The keypoint of T3 could be concluded that: in the whole training process, both the assistant task and the target task share the samples of the assistant task. At first, the used LLM  [^1] GPT-4o is recommended, while any other large models with sufficient parameter capacity to possess emergence could be considered. acquires experience in handling the assistant task through the labeled data, which means a deeper understanding of the samples. Subsequently, LLM iteratively tries to accomplish the target task on these same samples and gain the experience for the target task by leveraging the previous learned contextual experiences, as well as being modulated by the predefined conditions. Finally, the model applies the experiences required for both the assistant and target tasks during the test process. Accordingly, a general workflow for task-agnostic T3 is shown in Figure 3 from Appendix A.1.
4 Experiments
4.1 Settings
4.1.1 Datasets
4.1.2 Evaluation Metrics
Factscore We use FactSumm and ChatGPT for Factscore calculation. FactSumm uses named-entity recognition and relation extraction to extract facts from the source text and the generated summary to evaluate fact consistency by comparing extracted facts represented as triples (subject, relation, object) Zhang et al. (2021). ChatGPT evaluates factual inconsistency in summaries, outperforming previous methods with fewer hallucinations Luo et al. (2023).
4.1.3 Baseline Models
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
BBC summary
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o* 40.19 | 41.71 18.14 | 21.45 36.97 | 39.44 10.86 | 10.94 63.64 | 70.12
GPT-4-turbo 37.33 | 36.83 14.76 | 14.14 33.73 | 33.27 8.51 | 7.69 60.91 | 69.01
GPT-3.5-turbo* 39.36 | 37.88 17.75 | 16.95 36.04 | 34.78 10.50 | 8.58 57.52 | 47.80
Gemini-1.5-pro* 33.62 | 35.10 11.48 | 12.58 30.36 | 31.69 5.05 | 6.47 25.88 | 60.49
Gemini-1.0-pro* 34.90 | 33.58 12.80 | 13.92 31.75 | 31.17 5.99 | 7.22 35.76 | 55.61
Claude-3.5-sonnet* 39.08 | 43.54 16.59 | 21.13 35.88 | 40.36 9.70 | 13.00 63.43 | 65.83
Claude-3.0-sonnet 41.23 | 40.90 18.26 | 18.08 37.78 | 37.39 11.08 | 10.74 60.02 | 50.43
Average 37.96 | 38.51 15.68 | 16.89 34.64 | 35.44 8.81 | 9.23 52.45 | 59.90
NarraSum
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o* 31.14 | 28.86 7.31 | 8.58 28.44 | 25.66 2.68 | 2.55 97.75 | 99.69
GPT-4-turbo* 29.14 | 27.27 7.72 | 7.42 25.44 | 24.45 3.36 | 3.60 97.75 | 97.44
GPT-3.5-turbo* 28.60 | 25.28 7.52 | 5.88 26.50 | 23.00 2.79 | 2.15 95.20 | 95.91
Gemini-1.5-pro* 25.99 | 26.95 5.21 | 5.39 25.00 | 25.32 1.75 | 1.99 93.26 | 93.57
Gemini-1.0-pro 28.97 | 28.25 6.26 | 5.74 25.45 | 26.36 1.65 | 2.09 92.85 | 91.22
Claude-3.5-sonnet 29.19 | 29.78 7.17 | 7.79 26.73 | 27.23 3.10 | 3.23 97.04 | 98.14
Claude-3.0-sonnet* 29.90 | 25.87 6.75 | 5.47 26.45 | 22.51 2.74 | 2.05 94.69 | 95.61
Average 28.99 | 27.47 6.85 | 6.61 26.29 | 24.93 2.58 | 2.52 95.51 | 95.94
NLQuAD
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o* 24.37 | 31.00 8.48 | 13.03 22.71 | 29.27 0.33 | 1.42 59.81 | 63.95
GPT-4-turbo* 24.19 | 23.66 7.39 | 7.36 22.16 | 21.62 0.44 | 0.36 53.33 | 49.16
GPT-3.5-turbo* 21.50 | 22.21 7.75 | 8.30 20.07 | 20.69 0.27 | 0.27 55.30 | 44.64
Gemini-1.5-pro* 22.99 | 24.19 6.97 | 7.97 21.34 | 22.44 0.25 | 0.47 50.85 | 65.09
Gemini-1.0-pro* 26.26 | 28.11 9.29 | 11.05 24.55 | 26.48 0.54 | 1.25 64.57 | 71.64
Claude-3.5-sonnet* 32.03 | 40.17 12.80 | 18.74 30.06 | 37.92 2.35 | 5.00 65.71 | 82.76
Claude-3.0-sonnet* 32.35 | 33.72 12.70 | 13.27 30.16 | 31.22 1.97 | 2.37 58.89 | 61.69
Average 26.24 | 29.01 9.34 | 11.39 24.44 | 27.09 0.88 | 1.59 58.35 | 62.70
FairytaleQA
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o* 13.34 | 14.86 3.15 | 3.72 12.34 | 13.78 0.04 | 0.07 98.55 | 98.92
GPT-4-turbo* 14.68 | 14.14 3.51 | 3.36 13.45 | 12.84 0.09 | 0.10 97.44 | 97.67
GPT-3.5-turbo 12.22 | 12.23 2.98 | 3.07 11.31 | 10.74 0.03 | 0.04 99.82 | 97.37
Gemini-1.5-pro* 15.40 | 17.56 3.80 | 4.34 14.30 | 15.46 0.10 | 0.13 97.23 | 98.53
Gemini-1.0-pro* 14.14 | 13.43 3.25 | 2.43 14.22 | 13.39 0.03 | 0.14 97.84 | 94.24
Claude-3.5-sonnet* 15.97 | 20.96 3.73 | 6.07 14.94 | 19.52 0.17 | 0.47 99.56 | 99.97
Claude-3.0-sonnet* 17.20 | 18.91 4.08 | 4.90 15.81 | 17.35 0.17 | 0.33 97.47 | 98.92
Average 14.71 | 16.01 3.50 | 3.98 13.77 | 14.73 0.09 | 0.18 98.27 | 97.95
4.2 Implementation Details
The baseline summary of each text for each LLM is generated by using original text and the following prompt: "You are one helpful text assistant. Based on the input text, provide a summary in proper length. Aim to maximize Performance scores". [^2] We add this to ensure fairness, since T3 use these words in its training process, according to Appendix A.3. Correspondently, each LLM with T3 is guided by the experience and prompts derived from training process.
For BBC summary and NarraSum, the provided summary was used as the reference, while for FairytaleQA and NLQuAD datasets, the original text was used as the reference. As mentioned in Part 3.4, T3 can also deal with QA datasets for summarization task. When the dataset does not provide reference summaries, the generated summary can be directly compared to the original text through certain automated metrics tools (e.g., ROUGE) Louis and Nenkova (2009). According to Saggion and Poibeau (2013), the overlap of the generated summaries with the original text is important in assessing information coverage and the validity.
4.3 Results
Table 1 illustrates the performance of all used models with (right "T3" column) and without (left "w/o" column) T3 on all adopted datasets. The better one of the two values from each row of the two columns is marked in bold. Especially, for each row where all results from columns with T3 are better than those columns without T3, the row is colored blue. Models corresponding to results with significant differences between w/o and T3 are superscripted with *, where the detailed calculation process is given in Appendix A.12. From this table, we observe that:
(1) LLMs using T3 outperforms themselves without T3 under most metrics of most datasets, which demonstates the feasibility and effectiveness of T3. Specially, the average improvement rates of GPT-4o, Gemini-1.5-pro, and Claude-3.5-sonnet using T3 compared to the baselines without T3 across different tasks are calculated and shown that these LLMs with T3 nearly improves the ROUGE score by 14%, the BLEU score by 35%, and the Factscore score by 16%. Most of the comparison groups (82.14%) showed significant differences with significance_level=0.05 and null_hypothesis that there is no significant difference between the means of the two groups.
4.4 Ablation Study
We further analyze the influence of different components of T3.
Table LABEL:tab:Table3 in Appendix A.11 shows the experimental results for all used models after removing the summary experience component $Exp_{Sum}$ or the QA generation experience component $Exp_{QA}$, on all used datasets. It could be found that: (1) For models with less reasoning ability, longer inputs and complex instructions might diminish their performance mainly in their $Exp_{Sum}$ component than their $Exp_{QA}$ component, while for more advanced models, both the two components could be utilized to enhance their overall effectiveness. (2) Since $Exp_{Sum}$ and $Exp_{QA}$ is run for the target task and the assistant task, respectively, it might be demonstrated the importance of the assistant task in the whole T3 workflow.
5 Conclusion
Limitations
Ethics Statement
Appendix A Appendix
A.1 General workflow of task-agnostic T3
The figure below shows the general task-agnostic workflow of T3, i.e., the overall workflow of T3 when both the assistant and target tasks are not identified. The workflow facilitates the matching of different assistant and target tasks in practice.
Figure 3: General workflow of task-agnostic T3
@@ A.2 Prompt for QA Generation and Experience Updating in Training Process
You are a helpful text assistant skilled at learning new experiences. First, please help us with the following:
Based on the input text, provided questions, and QA pair generation experience, generate questions and corresponding answers that reflect all subjects and key points of the input article. The format of the generated QA pairs should match that of the provided question pairs.
Update the learning experience for generating better QA pairs during the question generation process by comparing the generated questions to the original question pairs based on the original text or other methods. Consider all key points needed to create good QA pairs, and ensure the learning experience is general. DO NOT INCLUDE DETAILED EXAMPLES BUT PROVIDE RULES AS DETAILED AS POSSIBLE. DO NOT EXCEED 600 WORDS.
@@
{ "Generated_QA_pairs": { "1": "Question": "xxxx", "Answer": "xxxx" , "2": "Question": "xxxx", "Answer": "xxxx" }, "QA_generation_experience": "Answer point by point text in one line" }